(w) IN A SURVEY OF 20 PATIENTS WHO SMOKED, THE FOLLOWING DATA WERE OBTAINED. EACH VALUE REPRESENTS THE NUMBER OF CIGARETTES THE PATIENT SMOKED PER DAY. CONSTRUCT A FREQUENCY DISTRIBUTION, USING SIX CLASSES.

(10,22,11,13,5) (8,13,9,12,11) (6,17,18,15,16) (14,19,14,15,11)
1
Expert's answer
2020-01-22T09:06:22-0500

Smallest value = 5

Largest value = 22

Range "=22-5=17"

Number of classes =6

Class width "=\frac{17}{6}=3" (round off to the nearest whole number)